OpenStudy (anonymous):

g(x) = -2x2 - 9 Find: g(3)

jimthompson5910 (jim_thompson5910):

replace every x with 3 to get g(x) = -2x^2 - 9 g(3) = -2(3)^2 - 9 g(3) = ???

OpenStudy (anonymous):

So, would that make it -27?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thank you.!

jimthompson5910 (jim_thompson5910):

sure thing